# Electronic modes induced by spin and charge perturbations in Mott and Kondo insulators

### v1

@@ Abstract
Electronic band structures usually remain unaffected by doping via a chemical-potential shift or by increasing the temperature in conventional band insulators. In contrast, it has been shown that those of Mott and Kondo insulators can change by doping or by increasing the temperature: electronic modes are induced within the band gap, exhibiting momentum-shifted magnetic dispersion relations from the band edges. Here, by generalizing the underlying mechanism of the remarkable strong-correlation effects, this study shows that various spin and charge perturbations, including magnetization of spin-gapped Mott and Kondo insulators, can induce electronic modes. The emergent modes can change the band structure, provided that a macroscopic number of spins or charges are excited by the perturbations at a moment. The origins and dispersion relations of the emergent modes are elucidated by investigating the selection rules and using the Bethe ansatz and effective theory for weak inter-unit-cell hopping. The theoretical results are validated by numerical calculations for the one- and two-dimensional Hubbard models, periodic Anderson models, Kondo lattice models, and ladder and bilayer Hubbard models. This study significantly advances the fundamental understanding of the changes in the band structure driven by spin and charge perturbations and paves the way for band-structure engineering in strong-correlation electronics, utilizing the unconventional characteristics of strongly correlated insulators.
@@ I Introduction
The outline of this paper is as follows: In Sec. II, the models, parameters, and quantities considered in this paper are defined. The details of the methods are presented. In Sec. III, the spectral properties of unperturbed systems are reviewed for comparison with those of perturbed systems shown in later sections. In Sec. IV, the spectral properties of charge-perturbed systems are considered. After reviewing how and why electronic states are induced by doping Kohno1DHub ; Kohno2DHub ; KohnoDIS ; KohnoRPP ; KohnoHubLadder ; KohnoKLM ; KohnoGW ; KohnoAF ; KohnoSpin ; Kohno1DtJ ; Kohno2DtJ ; KohnoMottT ; KohnoTinduced , the characteristics are extended to those of the emergent electronic states in undoped charge-perturbed systems with nonzero values of the charge ($\eta$) quantum number, using relations based on $\eta$-SU(2) symmetry. Remarks on nonzero-$\eta$ states at half filling are made, because photoinduced $\eta$-pairing states have recently attracted considerable attention entropyCoolingDMFT ; etaPairingSuper ; nonthermalSuper ; KanekoEtaPairing ; etaPairingHubLad , which are generally non-superconducting states. The dispersion relations of electronic modes induced by charge fluctuation are also derived. In Sec. V, the spectral properties of spin-perturbed systems are considered. The origins and dispersion relations of magnetization-induced electronic modes in spin-gapped Mott and Kondo insulators are explained. Based on the results, the emergent electronic modes in unmagnetized spin-perturbed systems with nonzero values of the spin quantum number are explained. The dispersion relations of electronic modes induced by spin fluctuation are also derived. In Sec. VI, the differences in spectral features between the strongly correlated insulators and conventional band insulators are discussed to elucidate the strong-correlation effects. The relevance of spin-charge separation in strongly correlated insulators to the doping-induced modes and magnetization-induced modes is clarified. Implications of this study are described. The paper is summarized in Sec. VII.
The theoretical and numerical results indicate that the spectral weights of the electronic modes induced by spin and charge perturbations can become comparable to those of the conventional bands, provided that a macroscopic number of spins or charges are excited by the perturbations at a moment; the spectral weights of the induced modes are proportional to the density of the excited spins or charges in the small-density regime. The dispersion relations of the emergent electronic modes can be expressed as combinations of the electronic dispersion relation and spin or charge dispersion relation of unperturbed systems. These results can be applied to the basic principle of band-structure engineering for strong-correlation electronics, thus guiding the tuning and selection of perturbations and strongly correlated materials.
@@ II.1 Models
where $\langle i,j\rangle$ means that sites $i$ and $j$ are nearest neighbors. Here, $c^{\dagger}_{i,\sigma}$ and $n^{c}_{i,\sigma}$ ($f^{\dagger}_{i,\sigma}$ and $n^{f}_{i,\sigma}$) denote the creation and number operators of a conduction orbital (localized orbital) electron with spin $\sigma$ for $\sigma=\uparrow,\downarrow$ at a site $i$, respectively, and ${\bm{S}}^{c}_{i}$ (${\bm{S}}^{f}_{i}$) denotes the spin operator of a conduction orbital (localized orbital) electron at a site $i$; ${\bm{S}}^{\lambda}_{i}=(S^{\lambda,x}_{i},S^{\lambda,y}_{i},S^{\lambda,z}_{i})$ for $\lambda=c$ and $f$. In the KLM, each localized orbital has one electron. Hereinafter, $t_{i,j}=t$ in chains and square-lattice planes, and $t_{i,j}=t_{\perp}$ between chains for the ladder HM and between square-lattice planes for the bilayer HM.
In this paper, the ground state of the particle–hole symmetric system ($\Delta=\frac{U}{2}$, $\mu=0$, and $H=0$) is considered the unperturbed state. The PAM for $U=2\Delta$ is called the symmetric PAM. The parameters $t,t_{\perp},t_{\rm K},U$, $\Delta$, and $J_{\rm K}$ are assumed to be positive unless otherwise mentioned. In a doped system (magnetized system), the chemical potential $\mu$ (magnetic field $H$) is adjusted such that the ground state has the corresponding doping concentration (magnetization) unless otherwise mentioned.
The numbers of electrons, sites, and unit cells are denoted by $N_{\rm e}$, $N_{\rm s}$, and $N_{\rm u}$, respectively. A unit cell is a site in the 1D and 2D HMs ($N_{\rm u}=N_{\rm s}$), a pair of sites connected by $t_{\perp}$ in the ladder and bilayer HMs ($N_{\rm u}=\frac{N_{\rm s}}{2}$), and a site with a conduction orbital and localized orbital in the PAM and KLM ($N_{\rm u}=N_{\rm s}$). The number of doped holes from half filling is denoted by $N_{\rm h}$: $N_{\rm h}=N_{\rm s}-N_{\rm e}$ in the HM and $2N_{\rm u}-N_{\rm e}$ in the PAM and KLM. The hole-doping concentration $\delta$ is defined as $\delta=\frac{N_{\rm h}}{N_{\rm s}}(=1-\frac{N_{\rm e}}{N_{\rm s}})$ in the HM and $\delta=\frac{N_{\rm h}}{2N_{\rm u}}(=1-\frac{N_{\rm e}}{2N_{\rm u}})$ in the PAM and KLM. At half filling, $\delta=N_{\rm h}=0$. In the hole-doped case, $\delta>0$ and $N_{\rm h}>0$; in the electron-doped case, $\delta<0$ and $N_{\rm h}<0$. Hereinafter, the number of electrons at half filling is assumed to be even, and the unit of $\hbar=1$ is used. The total spin quantum number and $z$ component of the spin are denoted as $S$ and $S^{z}$, respectively. The maximum value of $S$ at half filling is denoted as $S_{\rm max}$: $S_{\rm max}=\frac{N_{\rm s}}{2}$ in the HM and $N_{\rm u}$ in the PAM and KLM. The electronic spin opposite to $\sigma$ is denoted as $\bar{\sigma}$}: $\bar{\sigma}$$\downarrow$ and $\uparrow$ for $\sigma=\uparrow$ and $\downarrow$, respectively. The $z$ component of the spin of an electron is denoted as $s^{z}$: $s^{z}=\frac{1}{2}$ and $\frac{1}{2}$ for $\sigma=\uparrow$ and $\downarrow$, respectively.
@@ II.2 Spectral function and dynamical structure factors
The spectral function of a state $|X\rangle$ with energy $E_{X}$ at zero temperature is defined as
@@
$A^{\lambda,\sigma}_{X}({\bm{k}},\omega)$ $=\frac{1}{Z_{X}}\sum_{n}\left[\begin{array}[]{r}|\langle n|\lambda^{\dagger}_{{\bm{k}},\sigma}|X\rangle|^{2}\delta(\omega-E_{n}+E_{X})\\ +|\langle n|\lambda_{{\bm{k}},\sigma}|X\rangle|^{2}\delta(\omega+E_{n}-E_{X})\end{array}\right]$ (10)
with $Z_{X}=\langle X|X\rangle$. Here, $|n\rangle$ denotes the $n$th eigenstate of the Hamiltonian with energy $E_{n}$, and $\lambda^{\dagger}_{{\bm{k}},\sigma}$ denotes the Fourier transform of $\lambda^{\dagger}_{i,\sigma}$ for $\lambda=c$ and $f$. In the HM and KLM, $A^{f,\sigma}_{X}({\bm{k}},\omega)=0$. The eigenstates of the Hamiltonian are assumed to be normalized to one.
@@
in the PAM and KLM, where $e_{m}$ denotes the excitation energy of $|m\rangle$ from the ground state $|{\rm GS}\rangle$, and $S^{\lambda,\alpha}_{\bm{k}}$ denotes the Fourier transform of $S^{\lambda,\alpha}_{i}$ for $\lambda=c$ and $f$; $S^{\lambda,\alpha}_{i}$ denotes the $\alpha$-component of ${\bm{S}}^{\lambda}_{i}$ for $\alpha=x$, $y$, and $z$. The charge fluctuation of a momentum $\bm{k}$} is caused by the following operator:
@@
The ground state in the subspace of $N_{\rm h}=n$ is denoted as $|{\rm GS}\rangle^{N_{\rm h}=n}$ regardless of the value of $\mu$ for $H=0$. In this paper, unless otherwise mentioned, the ground state means the ground state at half filling for $H=0$, $|{\rm GS}\rangle^{N_{\rm h}=0}$, which is denoted by $|{\rm GS}\rangle$. The energies of $|{\rm GS}\rangle$ and $|{\rm GS}\rangle^{N_{\rm h}=n}$ are denoted as $E_{\rm GS}$ and $E_{\rm GS}^{N_{\rm h}=n}$, respectively.
@@ II.3 Methods
The spectral function of $|{\rm GS}\rangle^{N_{\rm h}=-n}$ for $n>0$ was obtained using that of $|{\rm GS}\rangle^{N_{\rm h}=n}$ via Eq. (169). The spectral function of $S^{c,z}_{\bm{q}}|{\rm GS}\rangle$ and ${\tilde{n}}^{c}_{\bm{q}}|{\rm GS}\rangle$ in the HM and that of $\frac{1}{\sqrt{2}}(S^{c,z}_{\bm{q}}-S^{f,z}_{\bm{q}})|{\rm GS}\rangle$ and $\frac{1}{\sqrt{2}}({\tilde{n}}^{c}_{\bm{q}}-{\tilde{n}}^{f}_{\bm{q}})|{\rm GS}\rangle$ in the PAM and KLM were calculated using the CPT, where $|X\rangle=\sum_{i}e^{i{\bm{q}}\cdot{\bm{r}}_{i}}O_{i}|{\rm GS}\rangle$ in Eq. (10) for $O_{i}=S^{c,z}_{i}$, ${\tilde{n}}^{c}_{i}$, $\frac{1}{\sqrt{2}}(S^{c,z}_{i}-S^{f,z}_{i})$, and $\frac{1}{\sqrt{2}}({\tilde{n}}^{c}_{i}-{\tilde{n}}^{f}_{i})$.
The local Green’s functions for the intermediate electron density, magnetization ($S^{z}$) density, spin ($S$) density, and $\eta$ density (Sec. III.1.2) were obtained as the weighted averages of the local Green’s functions of larger and smaller densities on a cluster in the CPT CPTHanke . The numerical calculations were performed for perturbed and unperturbed states with even numbers of electrons.
The numerical results with Gaussian broadening were obtained through deconvolution from those with Lorentzian broadening Kohno1DHub ; Kohno2DHub ; KohnoDIS ; KohnoRPP ; KohnoHubLadder ; KohnoKLM ; KohnoGW ; KohnoAF ; KohnoSpin ; Kohno1DtJ ; Kohno2DtJ ; KohnoMottT ; KohnoTinduced ; KohnottpHub ; KohnottpJ .

### v2

@@ Abstract
Electronic band structures usually remain unaffected by doping via a chemical-potential shift or by increasing the temperature in conventional band insulators. In contrast, it has been shown that those of Mott and Kondo insulators can be altered by doping or by increasing the temperature: electronic modes are induced within the band gap, exhibiting momentum-shifted magnetic dispersion relations from the band edges. Here, this study demonstrates that the underlying mechanism of the remarkable strong-correlation effects can be generalized to the emergence of electronic modes caused by various spin and charge perturbations, including magnetization of spin-gapped Mott and Kondo insulators. These emergent modes can alter the band structure if a macroscopic number of spins or charges are excited by the perturbations at a given moment. The origins and dispersion relations of these emergent modes, particularly why and how the dispersion relations depend on the momentum and energy of the perturbations, are elucidated by investigating the selection rules and using the Bethe ansatz and the effective theory for weak inter-unit-cell hopping. The validity and generality of the theoretical results across different models and spatial dimensions are verified by numerical calculations for the one- and two-dimensional Hubbard models, periodic Anderson models, Kondo lattice models, and ladder and bilayer Hubbard models. This study provides crucial insights into why and how spin and charge perturbations can alter the band structure in strongly correlated insulators, thereby paving the way for band-structure engineering in strong-correlation electronics, which enables previously unexplored functionalities by exploiting the unconventional characteristics revealed in this paper.
@@ I Introduction
The outline of this paper is as follows: In Sec. II, the models, parameters, and quantities considered in this paper are defined. The details of the methods are presented. In Sec. III, the spectral properties of unperturbed systems are reviewed for comparison with those of perturbed systems shown in later sections. In Sec. IV, the spectral properties of charge-perturbed systems are considered. After reviewing how and why electronic states are induced by doping Kohno1DHub ; Kohno2DHub ; KohnoDIS ; KohnoRPP ; KohnoHubLadder ; KohnoKLM ; KohnoGW ; KohnoAF ; KohnoSpin ; Kohno1DtJ ; Kohno2DtJ ; KohnoMottT ; KohnoTinduced , the characteristics are extended to those of the emergent electronic states in undoped charge-perturbed systems with nonzero values of the charge ($\eta$) quantum number, using relations based on $\eta$-SU(2) symmetry. Remarks on nonzero-$\eta$ states at half filling are made, because photoinduced $\eta$-pairing states have recently attracted considerable attention entropyCoolingDMFT ; etaPairingSuper ; nonthermalSuper ; KanekoEtaPairing ; etaPairingHubLad , which are generally nonsuperconducting states. The dispersion relations of electronic modes induced by charge fluctuation are also derived. In Sec. V, the spectral properties of spin-perturbed systems are considered. The origins and dispersion relations of magnetization-induced electronic modes in spin-gapped Mott and Kondo insulators are explained. Based on the results, the emergent electronic modes in unmagnetized spin-perturbed systems with nonzero values of the spin quantum number are explained. The dispersion relations of electronic modes induced by spin fluctuation are also derived. In Sec. VI, the differences in spectral features between the strongly correlated insulators and conventional band insulators are discussed to elucidate the strong-correlation effects. The relevance of spin-charge separation in strongly correlated insulators to the doping-induced modes and magnetization-induced modes is clarified. Implications of this study are described. The paper is summarized in Sec. VII.
The theoretical and numerical results indicate that the spectral weights of the electronic modes induced by spin and charge perturbations can become comparable to those of the conventional bands, provided that a macroscopic number of spins or charges are excited by the perturbations at a given moment; the spectral weights of the induced modes are proportional to the density of the excited spins or charges in the small-density regime. The dispersion relations of the emergent electronic modes can be expressed as combinations of the electronic and either the spin or the charge dispersion relations of unperturbed systems. These results can be applied to the basic principle of band-structure engineering for strong-correlation electronics, thus guiding the tuning and selection of perturbations and strongly correlated materials.
@@ II.1 Models
where $\langle i,j\rangle$ means that sites $i$ and $j$ are nearest neighbors. Here, $c^{\dagger}_{i,\sigma}$ and $n^{c}_{i,\sigma}$ ($f^{\dagger}_{i,\sigma}$ and $n^{f}_{i,\sigma}$) denote the creation and number operators of a conduction orbital (localized orbital) electron with spin $\sigma$ for $\sigma=\uparrow,\downarrow$ at a site $i$, respectively, and ${\bm{S}}^{c}_{i}$ (${\bm{S}}^{f}_{i}$) denotes the spin operator of a conduction orbital (localized orbital) electron at a site $i$; ${\bm{S}}^{\lambda}_{i}=(S^{\lambda,x}_{i},S^{\lambda,y}_{i},S^{\lambda,z}_{i})$ for $\lambda=c$ and $f$. In the KLM, each localized orbital has one electron. Hereinafter, we set $t_{i,j}=t$ in chains and square-lattice planes, $t_{i,j}=t_{\perp}$ between chains for the ladder HM and between square-lattice planes for the bilayer HM, and $\Delta=\frac{U}{2}$ for the PAM; the PAM with $U=2\Delta$ is called the symmetric PAM. The parameters $t,t_{\perp},t_{\rm K},U$, $\Delta$, and $J_{\rm K}$ are assumed to be positive unless otherwise mentioned.
In this paper, the ground state at $\mu=H=0$ is considered the unperturbed state. In a doped system (magnetized system), the chemical potential $\mu$ (magnetic field $H$) is adjusted such that the ground state has the corresponding doping concentration (magnetization) unless otherwise mentioned.
The numbers of electrons, sites, and unit cells are denoted by $N_{\rm e}$, $N_{\rm s}$, and $N_{\rm u}$, respectively. A unit cell is a site in the 1D and 2D HMs ($N_{\rm u}=N_{\rm s}$), a pair of sites connected by $t_{\perp}$ in the ladder and bilayer HMs ($N_{\rm u}=\frac{N_{\rm s}}{2}$), and a site with a conduction orbital and localized orbital in the PAM and KLM ($N_{\rm u}=N_{\rm s}$). The number of doped holes from half filling is denoted by $N_{\rm h}$: $N_{\rm h}=N_{\rm s}-N_{\rm e}$ in the HM and $2N_{\rm u}-N_{\rm e}$ in the PAM and KLM. The hole-doping concentration $\delta$ is defined as $\delta=\frac{N_{\rm h}}{N_{\rm s}}(=1-\frac{N_{\rm e}}{N_{\rm s}})$ in the HM and $\delta=\frac{N_{\rm h}}{2N_{\rm u}}(=1-\frac{N_{\rm e}}{2N_{\rm u}})$ in the PAM and KLM. At half filling, $\delta=N_{\rm h}=0$. In the hole-doped case, $\delta>0$ and $N_{\rm h}>0$; in the electron-doped case, $\delta<0$ and $N_{\rm h}<0$. Hereinafter, the number of electrons at half filling is assumed to be even, and the unit of $\hbar=1$ is used. The total spin quantum number and $z$ component of the spin are denoted by $S$ and $S^{z}$, respectively. The maximum value of $S$ at half filling is denoted by $S_{\rm max}$: $S_{\rm max}=\frac{N_{\rm s}}{2}$ in the HM and $N_{\rm u}$ in the PAM and KLM. The electronic spin opposite to $\sigma$ is denoted by $\bar{\sigma}$}: $\bar{\sigma}$$\downarrow$ and $\uparrow$ for $\sigma=\uparrow$ and $\downarrow$, respectively. The $z$ component of the spin of an electron is denoted by $s^{z}$: $s^{z}=\frac{1}{2}$ and $\frac{1}{2}$ for $\sigma=\uparrow$ and $\downarrow$, respectively.
@@ II.2 Spectral function and dynamical structure factors
The spectral function of a state $|X\rangle$ at zero temperature is defined as
@@
$A^{\lambda,\sigma}({\bm{k}},\omega)$ $=-\frac{1}{\pi}{\rm Im}G^{\lambda}_{{\bm{k}},\sigma}(\omega)$ (8)
using the following Green’s function:
$G^{\lambda}_{{\bm{k}},\sigma}(\omega)$ $=-\frac{i}{Z_{X}}\int_{0}^{\infty}dt{\rm e}^{i\omega t-\epsilon t}\langle X|\{\lambda_{{\bm{k}},\sigma}(t),\lambda^{\dagger}_{{\bm{k}},\sigma}\}|X\rangle,$ (9)
$Z_{X}$ $=\langle X|X\rangle,\quad\lambda_{{\bm{k}},\sigma}(t)={\rm e}^{i{\cal H}t}\lambda_{{\bm{k}},\sigma}{\rm e}^{-i{\cal H}t}$
[…]
$A^{\lambda,\sigma}_{X}({\bm{k}},\omega)$ $=\frac{1}{Z_{X}}\sum_{n}\left[\begin{array}[]{r}|\langle n|\lambda^{\dagger}_{{\bm{k}},\sigma}|X\rangle|^{2}\delta(\omega-E_{n}+E_{X})\\ +|\langle n|\lambda_{{\bm{k}},\sigma}|X\rangle|^{2}\delta(\omega+E_{n}-E_{X})\end{array}\right],$ (12)
where $|n\rangle$ denotes the $n$th eigenstate of the Hamiltonian with energy $E_{n}$. The eigenstates of the Hamiltonian are assumed to be normalized to unity. In the HM and KLM, $A^{f,\sigma}_{X}({\bm{k}},\omega)=0$.
@@
in the PAM and KLM, where $e_{m}$ denotes the excitation energy of $|m\rangle$ from the ground state $|{\rm GS}\rangle$, and $S^{\lambda,\alpha}_{\bm{k}}$ denotes the Fourier transform of $S^{\lambda,\alpha}_{i}$ for $\lambda=c$ and $f$; $S^{\lambda,\alpha}_{i}$ denotes the $\alpha$ component of ${\bm{S}}^{\lambda}_{i}$ for $\alpha=x$, $y$, and $z$. The charge fluctuation of a momentum $\bm{k}$} is caused by the following operator:
@@
The ground state in the subspace of $N_{\rm h}=n$ is denoted by $|{\rm GS}\rangle^{N_{\rm h}=n}$ regardless of the value of $\mu$ for $H=0$. In this paper, unless otherwise mentioned, the ground state means the ground state at half filling for $H=0$, $|{\rm GS}\rangle^{N_{\rm h}=0}$, which is denoted by $|{\rm GS}\rangle$. The energies of $|{\rm GS}\rangle$ and $|{\rm GS}\rangle^{N_{\rm h}=n}$ are denoted by $E_{\rm GS}$ and $E_{\rm GS}^{N_{\rm h}=n}$, respectively.
@@ II.3 Methods
The spectral function of $|{\rm GS}\rangle^{N_{\rm h}=-n}$ for $n>0$ was obtained using that of $|{\rm GS}\rangle^{N_{\rm h}=n}$ via Eq. (171). The spectral functions of magnetized states in the HM and PAM were obtained using the Shiba transformation [Eqs. (112) and (116); Sec. V.5]. The spectral functions of $S^{c,z}_{\bm{q}}|{\rm GS}\rangle$ and ${\tilde{n}}^{c}_{\bm{q}}|{\rm GS}\rangle$ in the HM and those of $\frac{1}{\sqrt{2}}(S^{c,z}_{\bm{q}}-S^{f,z}_{\bm{q}})|{\rm GS}\rangle$ and $\frac{1}{\sqrt{2}}({\tilde{n}}^{c}_{\bm{q}}-{\tilde{n}}^{f}_{\bm{q}})|{\rm GS}\rangle$ in the PAM and KLM were calculated using CPT on sixteen-site clusters in the 1D, 2D, and ladder HMs, six-unit-cell clusters in the bilayer HM and 2D PAM, eight-unit-cell clusters in the 1D PAM, and nine-unit-cell clusters in the 1D and 2D KLMs. These calculations were performed using Eq. (12), where $|X\rangle$ is represented by $\sum_{i}e^{i{\bm{q}}\cdot{\bm{r}}_{i}}O_{i}|{\rm GS}\rangle$ for $O_{i}=S^{c,z}_{i}$, ${\tilde{n}}^{c}_{i}$, $\frac{1}{\sqrt{2}}(S^{c,z}_{i}-S^{f,z}_{i})$, and $\frac{1}{\sqrt{2}}({\tilde{n}}^{c}_{i}-{\tilde{n}}^{f}_{i})$, under the assumption that these states can be effectively identified as the dominant states in the dynamical structure factors.
The local Green’s functions for the intermediate electron density, magnetization ($S^{z}$) density, spin ($S$) density, and $\eta$ density (Sec. III.1.2) were obtained as the weighted averages of the local Green’s functions of larger and smaller densities on a cluster in CPT CPTHanke . The numerical calculations were performed for perturbed and unperturbed states with even numbers of electrons.
The results with Gaussian broadening were obtained through deconvolution from those with Lorentzian broadening Kohno1DHub ; Kohno2DHub ; KohnoDIS ; KohnoRPP ; KohnoHubLadder ; KohnoKLM ; KohnoGW ; KohnoAF ; KohnoSpin ; Kohno1DtJ ; Kohno2DtJ ; KohnoMottT ; KohnoTinduced ; KohnottpHub ; KohnottpJ .
In the figures presented in this paper, the numerical and theoretical results are shown for $U/t=6$ in the 1D HM, $U/t=9$ in the 2D HM, $U/t=8$ and $t_{\perp}/t=4$ in the ladder HM (except for Fig. 16 below), $U/t=16$ and $t_{\perp}/t=8$ in the bilayer HM, $U/t=2\Delta/t=14$ and $t_{\rm K}/t=4$ in the 1D PAM, $U/t=2\Delta/t=28$ and $t_{\rm K}/t=8$ in the 2D PAM, $J_{\rm K}/t=4$ in the 1D KLM, and $J_{\rm K}/t=8$ in the 2D KLM. For the spectral functions, the dynamical structure factors, and the imaginary part of the dynamical spin susceptibility, Gaussian broadening is used with the standard deviations of $0.1t$ for the 1D and 2D HMs, $0.16t$ for the ladder HM and 1D PAM, $0.32t$ for the bilayer HM and 2D PAM, $0.12t$ for the 1D KLM, and $0.24t$ for the 2D KLM.